homeschool12345:

Pleaseeeeee Helpppp I need help fast!!!!!!!!!!! Part A: The area of a square is (16x2 − 8x + 1) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. Part B: The area of a rectangle is (81x2 − 4y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work.

barstho:

Okay, so let's start by Part A: We have 16x^2-8x+1 To factor this, we need to split the middle term. So, let's find two factors that multiply to gives us 16 and add to give us -8. That would be -4 and -4. -4 x - 4 = 16 -4-4 = -8 So, we'll add that in for the middle term. 16x^2-4x-4x+1 Now we factor by grouping: (16x^2-4x)+(-4x+1) ==> Find the GCF for both groups 4x(4x-1)-1(4x-1) (4x-1)(4x-1) = (4x-1)^2 So, the length of the each side = 4x-1

jhonyy9:

@barstho i think you need getting here the real positiv root of this quadratic and this will be the length of this square

jhonyy9:

and what is about part B. ?

jhonyy9:

@barstho

jhonyy9:

in case of part B. us formula a^2 -b^2 = (a-b)(a+b)

barstho:

They aren't being responsive :\ So I don't want to just do it for them